OpenStudy (anonymous):

Help please?? A delicatessen offers 4 different breads, 4 cheeses, and 6 different meats. In how many ways can a sandwich be made with 1 bread, 2 cheese and 3 meats? 360 420 480 540

OpenStudy (kropot72):

Th required number of ways is given by: \[4\times 4C2\times6C3=you\ can\ calculate\]

OpenStudy (anonymous):

What do the C's stand for?

OpenStudy (kropot72):

\[\large nCr=\frac{n!}{r!(n-r)!}\] So 4C2 is \[\large 4C2=\frac{4\times3\times2\times1}{2\times1\times2\times1}=6\]

OpenStudy (kropot72):

And 6C3 is \[\large 6C3=\frac{6\times5\times4\times3\times2\times1}{3\times2\times1\times3\times2\times1}=20\]

OpenStudy (anonymous):

is it 480 then?

OpenStudy (kropot72):

Correct.